solve the incomplete quadratic equation by the square root method
1) x/4=9/x

Answer by RAY100(1637)

x/4 =9/x,,,,,cross multiply
.
x^2 =36
.
take square root of both sides
.
x= +/- 6
.
check,,,6/4 =9/6 = 3/2,,,ok
.
-6/4 =9/-6 = - 3/2,,,,ok